Hi so i can not figure out how to do this! i looked in my book and i cant find anything! so anyway i hope you can help! thank you:)




the sum of three consecutive odd integers is 465. What are the integers?

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

the sum of three consecutive odd integers is 465. What are the integers?
-------------
1st odd: 2x-1
2nd odd: 2x+1
3rd odd: 2x+3
------------------
Equation:
sum = 465
6x+3 = 465
6x = 462
x = 77
---
1st: 2x-1 = 153
2nd: 2x+1 = 155
3rd: 2x+3 = 157
--------------------
Cheers,
Stan H.

Answer by Alan3354(69443)

465/3 = 155, the middle number.
--> 153, 155, 157